Curriculum Vitae

Scott A. Klasky May 29, 1996

Social Security Number: 185-60-5000
Home Address:  10926 Jollyville Rd. Apt.1124.

Austin, TX 78759

Phone(home):  (512) 338-9324
Email:  scott@einstein.ph.utexas.edu

Education

• University of Texas, Center for Relativity, Ph.D. Physics, Summer 1994.

• Drexel University, B.S. Physics, 1989

Post-Doctoral Fellowships

• University of Texas, Center for Relativity, 1994-1995

Publications

• “Multigrid support with the DAGH package: Specifications and Applications” (w/ M. Choptuik et al.), Site report, 1995 .

• “A Technique for Tracking Apparent Horizons,” (w/ M. Huq et al.), to be submitted .

• “A Parallel Implementation of Multi-Grid in one dimension” (w/ R. Guenther), ( Site report, 1994).

• “Visualizing Complex Patterns in the Spread of Head and Neck Cancers,” (w/ L. Gray et al. ), The International Journal of Supercomputer Applications i7, 167 (1993).

• “Three-dimensional initial data for the collision of two black holes,” (w/ G. Cook et al. ), Physical Review D47, 1471 (1993).

• “Properties of gravitational “solitons” ” (w/ J. Centrella et al. ), Physical Review D43, 379 (1991).

Scott A. Klasky May 29, 1996

Awards and Distinctions • Member of Phi Beta Kappa, honor society, 1991

• Senior Honorable Mention Award at Drexel University, 1989

• Member and President of Sigma Pi Sigma at Drexel University, 1987 Professional Experience

• The University of Texas System Center for High Performance Computing, Department of Scientific Visualization, Research Assistant, 1989-1993. Developed Marching Cubes Algorithms and visualization of medical data.

• Drexel University, Relativity Group, Research Assistant, 1989. Coauthored a Computational/Theoretical paper on solitons on a Kasner background.

• Princeton Plasma Physics Laboratory, Diagnostics Division, Junior Physicist, 1987,1988. Developed software to analyze spectroscopy data. Developed hardware for use of monitoring the spectroscopy equipment.

• Dyna East Corporation, Junior Engineer, 1986. Simulated warheads

using Finite Element codes. Computer Languages and Other Skills

• FORTRAN, fluent; C, fluent; GL, fluent; DORE, fluent; Maple; Pascal.

Numerical Algorithm Experience • Parallel Multi Grid

• Multi Grid

• Adaptive Techniques

• Domain Decomposition

• Richardson extrapolation

• Deferred Correction

• Conjugate Gradient Methods

• Jacobi, Gauss Seidel, S.O.R.

Invited Professional Talks

• “Numerical techniques for tracking boosted black holes”, University of Texas, Relativity Seminar, 1995

• “Numerical techniques for the Initial Value Problem for Two Black Holes”, University of Texas, Final Defense, 1994

• “Multi Level Adaptive Tehniques, MLAT, for Elliptic Equations”, University of Illinois, NCSA, 1994

• “Multi Level Adaptive Techniques for the Initial Value Problem of Black Holes” , Penn State University, 1993

• “Generating Highly accurate Black Hole Data from low order techniques”, University of Texas, Relativity Seminar, 1993

• “A three dimensional line S.O.R. method for the solution of the black hole initial value problem”, University of Illinois, NCSA, 1992

• “A three dimensional line S.O.R. method for the solution of the black hole initial value problem”, University of Texas, Relativity Seminar, 1992

• “The annihilation of a cosmic string loop”, University of Texas, Qualifier, 1991

• “Cosmic Strings, the seeds for galaxies in the early universe”, University of Texas, Relativity Seminar, 1991

• “Cosmic Strings, the seeds for galaxies in the early universe”, The University of Texas System Center for High Performance Computing, 1991

References

• Richard A. Matzner, Center for Relativity, University of Texas at Austin

• Matthew W. Choptuik, Center for Relativity, University of Texas at Austin

• Manish Parashar, Center for Relativity, University of Texas at Austin